# A Minimal Substitution Basis for the Kalmar Elementary Functions

### v2

@@ Abstract
We show that the class of Kalmar elementary functions can be inductively generated from the addition, the integer remainder and the base-two exponentiation, hence improving previous results by Marchenkov and Mazzanti. We also prove that the substitution basis defined by these three operations is minimal. Furthermore, we discuss alternative substitution bases under arity constraints.
@@ 1 Introduction
Given an integer $r\geq 1$ and a $r$-ary operation $f$ on $\mathbb{N}$, we sometimes write $f$ as $(x_{1},\ldots,x_{r})\mapsto f(x_{1},\ldots,x_{r})$ or even just as $f(x_{1},\ldots,x_{r})$. For example, we sometimes denote the addition as $x+y$.
We also apply the conventions $0^{0}=1$ (see Mendelson [9, Proposition 3.16]) and $\left\lfloor x/0\right\rfloor=0$ (see Mazzanti [8, Section 2.1]), so, in particular, by applying the definition of the integer remainder, $x\bmod y:=x-y\left\lfloor x/y\right\rfloor$, we have that $x\bmod 1=0$ and $x\bmod 0=x$.
@@
Given a set $F$ of operations on $\mathbb{N}$, we denote by $\langle F\rangle$ the closure under substitution of the set $F\cup C\cup P$, where $C$ is the set of constant operations on $\mathbb{N}$ (which can be identified with non-negative integers) and $V$ is the set of projections on $\mathbb{N}$ (which can be identified with variables). And we say that $F$ is a substitution basis for a given set $G$ of operations on $\mathbb{N}$ if, and only if, $\langle F\rangle=G$ (cf. Mazzanti [8, Section 2.2]). For instance, if $S:=\langle x+y,\ x^{y}\rangle$, then $x^{z+2}\in S$: indeed, the projection $(x,y,z)\mapsto z$ and the constant operation $(x,y,z)\mapsto 2$ belong to $S$, so, because of the closure under substitution of $S$, the operation $(x,y,z)\mapsto x^{z+2}$ also belongs to $S$. Observe that some authors do not admit the constant operations as basic elements in a substitution basis (cf. Marchenkov [6, Introduction]), but, as we are already admitting any variable $x_{n}$, it is not unreasonable to also admit any constant $n$.
@@
The Kalmar elementary functions are the operations on $\mathbb{N}$ that can be computed in iterated exponential time (cf. Marchenkov [6, Introduction] and Oitavem [2, Introduction]). They form the class $\mathcal{E}^{3}$ in the Grzegorczyk hierarchy (see Grzegorczyk [1]), which encompasses most of the operations on $\mathbb{N}$ that frequently arise in mathematical contexts. Notable examples include classical number-theoretic functions such as the factorial function $n!$ (see Robinson [3] and Prunescu & Sauras-Altuzarra [11]), the $n$-th prime function $p_{n}$ (see Prunescu & Shunia [15]), Euler’s totient function $\varphi(n)$ (see Prunescu & Sauras-Altuzarra [12]), the greatest common divisor $\gcd(m,n)$ (see Prunescu & Shunia [14]), and the C-recursive integer sequences with rational coefficients (see Prunescu [10] and Prunescu & Sauras-Altuzarra [13]). In practical terms, a significant portion of the real-world computation can be faithfully modeled within the class of Kalmar elementary functions, since the corresponding code avoids unbounded recursion.
For a long time, it was hypothesized that a finite quantity of elementary operations should suffice to form a substitution basis for the class of Kalmar elementary functions. Concretely, it was conjectured that $\langle x+y,\ x\dotdiv y,\ xy,\ \left\lfloor x/y\right\rfloor,\ x^{y}\rangle=\mathcal{E}^{3}$ (see Marchenkov [6, Introduction]). Building upon the foundational contributions of Marchenkov [5], Jones & Matiyasevich [4], and Matiyasevich [7], Mazzanti [8] proved this conjecture. Slightly later, Marchenkov [6, Corollary 2] established that $\langle x+y,\ x\bmod y,\ x^{2},\ 2^{x}\rangle=\mathcal{E}^{3}$. 1 improves Marchenkov’s result by eliminating the squaring operation from the above substitution basis, that is to say, it asserts that $\langle x+y,\ x\bmod y,\ 2^{x}\rangle=\mathcal{E}^{3}$. And Section 3 is devoted to show that this substitution basis is minimal: if $F\subsetneqq\{x+y,x\bmod y,2^{x}\}$, then $\langle F\rangle\subsetneqq\mathcal{E}^{3}$.
@@
In addition, Section 4 analyzes why certain selected sets cannot be substitution bases for the class of Kalmar elementary functions. And, in Section 5, we present some finite substitution bases for the class of univariate Kalmar elementary functions that consist of unary operations only, and a substitution basis for the whole class of Kalmar elementary functions that consists of exactly one binary operation.
@@ 2 A substitution basis for the Kalmar elementary functions
We start by stating Theorem 1, which is one of the main results obtained by Marchenkov [6, Corollary 2], already mentioned in Section 1.
@@
and, by again applying that $0\leq 2^{x+y}+x<2^{x+y}+y$ if $x<y$, the statement follows. ∎
@@ 3 Minimality of the main substitution basis
Recall that the definition of substitution basis that we gave in Section 1 permits to use constants when generating functions from a given basis.
@@
If $t(x)=t_{1}(x)\bmod t_{2}(x)$, then we can take $B=\max(B_{1},B_{2})$: indeed, for every non-negative integer $a$, we have that:
@@
The proof goes by induction on the elements of $\langle x+y,\ 2^{x}\rangle$. The statement is obvious for the constants and the variable $x$. And, if $t_{1}(x)$ and $t_{2}(x)$ are strictly increasing elements of $\langle x+y,\ 2^{x}\rangle$, then $t_{1}(x)+t_{2}(x)$ and $2^{t_{1}(x)}$ are also strictly increasing. ∎
@@
The proof goes by induction on the elements of $\langle x+y,\ x\bmod y\rangle$. The statement is obvious for the constants and the variable $x$. And, if $t_{1}(x)$ and $t_{2}(x)$ are elements of $\langle x+y,\ x\bmod y\rangle$ such that $t_{1}(x)<A_{1}x+B_{1}$ and $t_{2}(x)<A_{2}x+B_{2}$ for some non-negative integers $A_{1}$, $A_{2}$, $B_{1}$ and $B_{2}$, then it is clear that $t_{1}(x)+t_{2}(x)<(A_{1}+A_{2})x+(B_{1}+B_{2})$ and $t_{1}(x)\bmod t_{2}(x)<A_{1}x+B_{1}$. ∎
@@ 4 Some sets which are not substitution bases
The proof is the same as in Theorem 9, after replacing $x+1$ with $x\dotdiv y$ in the set. ∎
@@
Finally, we state an open problem, 1.
@@ 5 Other finite substitution bases with low-arity operations
Cantor’s pairing function, which we denote by $x\mid y]$, is a bijection that maps each pair $(x,y)$ of non-negative integers into the non-negative integer $(x+y)(x+y+1)/2+x$ (see Prunescu & Sauras-Altuzarra [12, Subsection 4.5]). In addition, we consider in this section the unary operations $l(x)$ and $r(x)$ such that $l([x\mid y])=x$ and $r([x\mid y])=y$.
@@
Given any two operations $a$ and $b$ in $\langle B\rangle$ such that $a$ is unary and $b$ is binary, let us consider the unary operations $a^{\prime}(x)=\left[a(l(x))\mid r(x)\right]$ and $b^{\prime\prime}(x)=\left[b(l(x),l(r(x)))\mid r(r(x))\right]$.
Let $U$ be the union of the sets $\{l(x),r(x),d(x),u(x),v(x)\}$, $\{a^{\prime}(x):a(x)\textrm{ is a unary operation in }B\}$, and $\{b^{\prime\prime}(x):b(x)\textrm{ is a binary operation in }B\}$.
We claim that $U$ is a finite basis for the univariate Kalmar elementary functions. And it suffices to prove by induction that, for each unary operation $f(x)$ in $\langle B\rangle$, the operation $f^{\prime}(x)$ belongs to $\langle U\rangle$, because then we can express $f(x)$ as $l(f^{\prime}(d(x)))$.
@@
Finally, Theorem 11, which is due to Emil Jeřábek (pers. comm.), solves the initial question that was posed at the beginning of this section.
@@
Proof.
By applying Theorem 10, there is an integer $k$ and a substitution basis $B$ for the class of univariate Kalmar elementary functions that consists of exactly $k$ unary operations $f_{0}(x)$, $\ldots$, $f_{k-1}(x)$ such that $f_{0}(x)=2(x+1)$.
In addition, it is easy to prove that, if $i$ is a non-negative integer, then the $i$-th iteration of $f_{0}(x)$, denoted, as usual, as $f_{0}^{i}(x)$, is equal to $2^{i}x+2^{i+1}-2$. For instance, $f_{0}^{0}(x)=x=2^{0}x+2^{0+1}-2$, $f_{0}^{1}(x)=f_{0}(x)=2(x+1)=2^{1}x+2^{1+1}-2$, and $f_{0}^{2}(x)=f_{0}(f_{0}(x))=2(2(x+1)+1)=2^{2}x+2^{2+1}-2$.

### v4

@@ Abstract
We show that the class of Kalmar elementary functions can be inductively generated from the addition, the integer remainder, and the base-two exponentiation, hence improving previous results by Marchenkov and Mazzanti. We also prove that the substitution basis defined by these three operations is minimal. Furthermore, we discuss alternative substitution bases under arity constraints.
@@ 1 Introduction
Given a set $F$ of operations on $\mathbb{N}$, we denote by $\langle F\rangle$ the closure under substitution of the set $F\cup C\cup P$, where $C$ is the set of constant operations on $\mathbb{N}$ (which can be identified with non-negative integers) and $V$ is the set of projections on $\mathbb{N}$ (which can be identified with variables). And we say that $F$ is a substitution basis for a given set $G$ of operations on $\mathbb{N}$ if, and only if, $\langle F\rangle=G$ (cf. Mazzanti [6, Section 2.2]). For instance, if $S:=\langle x+y,\ x^{y}\rangle$, then $x^{z+2}\in S$: indeed, the projection $(x,y,z)\mapsto z$ and the constant operation $(x,y,z)\mapsto 2$ belong to $S$, so, because of the closure under substitution of $S$, the operation $(x,y,z)\mapsto x^{z+2}$ also belongs to $S$. Observe that some authors do not admit the constant operations as basic elements in a substitution basis (cf. Marchenkov [5, Introduction]), but, as we are already admitting any variable $x_{k}$, it is not unreasonable to also admit any constant $k$.
@@
The Kalmar elementary functions are the operations on $\mathbb{N}$ that can be computed in iterated exponential time (cf. Marchenkov [5, Introduction] and Oitavem [8, Introduction]). They form the class $\mathcal{E}^{3}$ in the Grzegorczyk hierarchy (see Grzegorczyk [2]), which encompasses most of the operations on $\mathbb{N}$ that frequently arise in mathematical contexts. Notable examples include classical number-theoretic functions such as the factorial function $n!$ (see Prunescu & Sauras-Altuzarra [10]), the sums of binomial coefficients (see Shunia & Sauras-Altuzarra [17, Theorem 3.1]), the $n$-th prime function $p_{n}$ (see Prunescu & Shunia [14, Theorem 10.1]), Euler’s totient function $\varphi(n)$ (see Prunescu & Sauras-Altuzarra [11, Theorem 4.12]), the greatest common divisor $\gcd(m,n)$ (see Prunescu & Shunia [13]), and any C-recursive integer sequence with rational coefficients (see Prunescu [9], Prunescu & Sauras-Altuzarra [12], and Prunescu & Shunia [15]). In practical terms, a significant portion of the real-world computation can be faithfully modeled within the class of Kalmar elementary functions, since the corresponding code avoids unbounded recursion.
In 1964, Rödding [16] (in German) showed that a finite quantity of elementary operations suffices to form a substitution basis for the class of Kalmar elementary functions, which marked the beginning of the search for substitution bases consisting of “simple” arithmetic operations. This search was considered concluded in 2002 (see Marchenkov [5, Introduction]), when Mazzanti [6] proved that
$\langle x+y,\ x\dotdiv y,\ xy,\ \left\lfloor x/y\right\rfloor,\ x^{y}\rangle=\mathcal{E}^{3}$ (1)
(long before Mazzanti obtained this result, it had already been incorrectly attributed to Marchenkov by Matiyasevich and Jones [3, Section 2], and later by Clote [1, Theorem 3.66]; however, what Marchenkov [4, Theorem 1] had actually demonstrated was the fact that
$\langle x+1,\left\lfloor x/y\right\rfloor,x^{y},\varphi(x,y)\rangle=\mathcal{E}^{3},$
where $\varphi(x,y)$ is equal to the least index $i$ such that $a_{i}=0$, if $x>1$ and $y=\sum_{i=0}^{\infty}a_{i}x^{i}$; and equal to zero otherwise). Slightly later, in 2007, Marchenkov [5, Corollary 2] established that
$\langle x+y,\ x\bmod y,\ x^{2},\ 2^{x}\rangle=\mathcal{E}^{3}.$
1 improves Marchenkov’s result by eliminating the squaring operation from the above substitution basis, that is to say, it asserts that
$\langle x+y,\ x\bmod y,\ 2^{x}\rangle=\mathcal{E}^{3}.$
@@
And Section 3 is devoted to show that this substitution basis is minimal: if $F\subsetneqq\{x+y,x\bmod y,2^{x}\}$, then $\langle F\rangle\subsetneqq\mathcal{E}^{3}$. In addition, Section 4 discusses why certain sets cannot serve as substitution bases for the class of Kalmar elementary functions. And, in Section 5, we present some finite substitution bases for the class of univariate Kalmar elementary functions that consist of unary operations only, and a substitution basis for the whole class of Kalmar elementary functions that consists of exactly one binary operation.
@@ 2 A substitution basis for the Kalmar elementary functions
We start by stating Theorem 1, which is one of the main results obtained by Marchenkov [5, Corollary 2], already mentioned in Section 1.
@@
and, by applying that $0\leq x-y<2^{x+y}+x$ if $x\geq y$, the statement follows. ∎
@@ 3 Minimality of the main substitution basis
Recall that the definition of substitution basis that we gave in Section 1 allows the usage of constants when generating functions from a given basis.
@@
If $t(x)=t_{1}(x)\bmod t_{2}(x)$, then we can set $B=\max(B_{1},B_{2})$: indeed, for every non-negative integer $a$, we have that:
@@
The proof goes by induction on the elements of $\langle x+y,\ 2^{x}\rangle$. The statement is obvious for the variable $x$. And, if $t_{1}(x)$ and $t_{2}(x)$ are strictly increasing elements of $\langle x+y,\ 2^{x}\rangle$, then $t_{1}(x)+t_{2}(x)$ and $2^{t_{1}(x)}$ are also strictly increasing. ∎
@@
The proof goes by induction on the elements of $\langle x+y,\ x\bmod y\rangle$. The statement is obvious for the constants and the variable $x$. And, if $t_{1}(x)$ and $t_{2}(x)$ are elements of $\langle x+y,\ x\bmod y\rangle$ such that $t_{1}(x)<A_{1}x+B_{1}$ and $t_{2}(x)<A_{2}x+B_{2}$ for some non-negative integers $A_{1}$, $A_{2}$, $B_{1}$, and $B_{2}$, then it is clear that $t_{1}(x)+t_{2}(x)<(A_{1}+A_{2})x+(B_{1}+B_{2})$ and $t_{1}(x)\bmod t_{2}(x)\leq t_{1}(x)<A_{1}x+B_{1}$. ∎
@@ 4 Some sets which are not substitution bases
The proof is the same as that of Theorem 9, but replacing $x+1$ with $x\dotdiv y$ in the set. ∎
@@
Finally, we state a research question, 1.
@@ 5 Other finite substitution bases with low-arity operations
Cantor’s pairing function, which we denote by $x\mid y]$, is a bijection that maps each pair $(x,y)$ of non-negative integers into the non-negative integer $(x+y)(x+y+1)/2+x$ (see Prunescu & Sauras-Altuzarra [11, Subsection 4.5]). In addition, we consider in this section the unary operations $l(x)$ and $r(x)$ such that $l([x\mid y])=x$ and $r([x\mid y])=y$.
@@
Given any two operations $a(x)$ and $b(x,y)$ in $\langle B\rangle$ such that $a(x)$ is unary and $b(x,y)$ is binary, let us consider the unary operations $a^{\prime}(x)=\left[a(l(x))\mid r(x)\right]$ and $b^{\prime\prime}(x)=\left[b(l(x),l(r(x)))\mid r(r(x))\right]$.
Let $U$ be the union of the sets $\{l(x),r(x),d(x),u(x),v(x)\}$, $\{a^{\prime}(x):a(x)\textrm{ is a unary operation in }B\}$, and $\{b^{\prime\prime}(x):b(x,y)\textrm{ is a binary operation in }B\}$.
We claim that $U$ is a finite basis for the univariate Kalmar elementary functions: indeed, it suffices to prove by induction that, for each unary operation $f(x)$ in $\langle B\rangle$, the operation $f^{\prime}(x)$ belongs to $\langle U\rangle$, because then we can express $f(x)$ as $l(f^{\prime}(d(x)))$.
@@
Finally, Theorem 11 solves the initial question that was posed at the beginning of this section.
@@
First proof (due to Emil Jeřábek, pers. comm.).
By applying Theorem 10, there is a positive integer $k$ and a substitution basis $B$ for the class of univariate Kalmar elementary functions that consists of exactly $k$ unary operations $f_{0}(x)$, $\ldots$, $f_{k-1}(x)$ such that $f_{0}(x)=2(x+1)$.
In addition, it is easy to prove that, if $i$ is a non-negative integer, then the $i$-th iteration of $f_{0}(x)$, denoted, as usual, as $f_{0}^{i}(x)$, is equal to $2^{i}x+2^{i+1}\dotdiv 2$. For instance, $f_{0}^{0}(x)=x=2^{0}x+2^{0+1}\dotdiv 2$, $f_{0}^{1}(x)=f_{0}(x)=2(x+1)=2^{1}x+2^{1+1}\dotdiv 2$, and $f_{0}^{2}(x)=f_{0}(f_{0}(x))=2(2(x+1)+1)=2^{2}x+2^{2+1}\dotdiv 2$.
@@
Second proof.
The identities
$\gcd(x,y)=\left\lfloor\dfrac{(2^{x^{2}y(y+1)}\dotdiv 2^{x^{2}y})(2^{x^{2}y^{2}}\dotdiv 1)}{(2^{x^{2}y}\dotdiv 1)(2^{xy^{2}}\dotdiv 1)2^{x^{2}y^{2}}}\right\rfloor\bmod 2^{xy}$
and
$\nu_{2}(x)=\left\lfloor\frac{\gcd(x,2^{x})^{x+1}\bmod(2^{x+1}\dotdiv 1)^{2}}{2^{x+1}\dotdiv 1}\right\rfloor$
for the greatest common divisor and dyadic valuation respectively, were obtained by Mazzanti [6, Lemma 3.3 and Lemma 3.4]; and a slightly simpler representation of the greatest common divisor,
$\gcd(x,y)=\left(\left\lfloor\dfrac{5^{xy(xy+x+y)}}{(5^{x^{2}y}\dotdiv 1)(5^{xy^{2}}\dotdiv 1)}\right\rfloor\bmod 5^{xy}\right)\dotdiv 1,$
was later obtained by Prunescu & Shunia [13, Theorem 4.1]. It is hence clear that $\gcd(x,y)$ and $\nu_{2}(x)$ are elements of $\langle x+y,\ x\dotdiv y,\ xy,\ \left\lfloor x/y\right\rfloor,\ x\bmod y,\ 2^{x}\rangle$.
Now, let $\psi(x)=\left\lfloor x/\gcd(x,2^{x})\right\rfloor$ (the odd part of $x$), $\mu(x)=\left\lfloor(\psi(x)\dotdiv 5)/2\right\rfloor$ and, for every $k\in\mathbb{N}$,
$X_{k}(x)=1\dotdiv\left((x\dotdiv k)+(k\dotdiv x)\right)=\begin{cases}0&\textup{ if }x\neq k,\\ 1&\textup{ if }x=k\end{cases}$
and
$X_{\geq k}(x)=1\dotdiv(k\dotdiv x)=\begin{cases}0&\textup{ if }x<k,\\ 1&\textup{ if }x\geq k.\end{cases}$
In addition, let
[…]
We have that
$g(x,y)=X_{0}(y)2^{x}+X_{1}(y)\left(2x+5\right)+X_{2}(y)\left(\nu_{2}(x)+\mu(x)\right)+X_{3}(y)\left(\nu_{2}(x)\bmod\mu(x)\right)+X_{\geq 4}(y)xy,$
so $\langle g(x,y)\rangle\subseteq\langle x+y,\ x\dotdiv y,\ xy,\ \left\lfloor x/y\right\rfloor,\ x\bmod y,\ 2^{x}\rangle=\mathcal{E}^{3}$ (see Eq. 1).
And, on the other hand, we have that
$2^{x}=g(x,0),$
$x+y=g(2^{x}(2y+5),2)=g(g(2^{x},2y+5),2)=g(g(g(x,0),g(y,1)),2),$
and
$x\bmod y=g(2^{x}(2y+5),3)=g(g(2^{x},2y+5),3)=g(g(g(x,0),g(y,1)),3),$
so, by applying 1, we get $\mathcal{E}^{3}=\langle x+y,\ x\bmod y,\ 2^{x}\rangle\subseteq\langle g(x,y)\rangle$. ∎